Define a variable, then write an eguation. Do not solve. For six consecutive weeks, Connie lost the same amount of weight. Six weeks ago she weighed 145 pounds. She now weighs 125 pounds. How many pounds did Connie loose each week?
Answer by stanbon(75887)

Define a variable, then write an eguation. Do not solve. For six consecutive weeks, Connie lost the same amount of weight. Six weeks ago she weighed 145 pounds. She now weighs 125 pounds. How many pounds did Connie loose each week?
------------
Let " number of weeks from now" "x".
Let "her weight" be "w".
You have two points (-6,145), (0,125)
Slope = [145-125]/-6 = -20/6 = -10/3
Equation:
w = (-10/3)x + 125
---------------
The slope indicates change in weight each week: slope = -10/3 lbs
========================
Cheers,
Stan H.